# A Debiased Estimator for the Mediation Functional in Ultra-High-Dimensional Setting in the Presence of Interaction Effects

### v2

@@ Appendix A Proofs
$T_{1}=\frac{1}{\sqrt{n(\sigma_{1}^{2}\|\tau_{1}\|_{2}^{2}+\sigma_{2}^{2}\|\tau_{2}\|_{2}^{2}\|\gamma_{1}\|_{2}^{2})}}\left(\sum_{\left\{i:A_{i}=1\right\}}\tau_{1,i}\epsilon_{i}+\sum_{\left\{i:A_{i}=0\right\}}\tau_{2,i}U_{i}^{\top}\gamma_{1}\right)\,,$
which follows by cancelling $\sqrt{n}$ from the definition of $\sigma_{n}$ in Equation (A.29). Since the noise is centered i.i.d. sub-Gaussian random variables (Assumption 5.5), we have conditional on $\mathcal{G}_{n}=\sigma(\mathcal{D}_{1},\mathbf{X},\mathbf{M}_{c},A_{1:n})$, $T_{1}$ has mean 0 and variance 1. Furthermore, it is easy to check that Lyapunov’s condition is also satisfied with $\delta=1$ as $\sqrt{n}\|\tau_{1}\|_{\infty}+\sqrt{n}\|\tau_{2}\|_{\infty}=o_{p}(1)$ and $\|\gamma_{1}\|_{2}$ is bounded. Therefore, an application of the conditional central limit theorem (Bulinski, (2017)) yields:
@@
$\sum_{\{i:A_{i}=0\}}\tau_{2,i}^{*2}=\sum_{\{i:A_{i}=0\}}\left(\frac{1}{n_{c}}\mu^{\top}\Sigma_{c}^{-1}(X_{c,i}-\mu_{c})\right)^{2}$ $=\frac{1}{n_{c}^{2}}\sum_{\{i:A_{i}=0\}}\left(\mu^{\top}\Sigma_{c}^{-1/2}P_{i}\right)^{2}$
$=\frac{1}{n_{c}^{2}}\sum_{\{i:A_{i}=0\}}\mu^{\top}\Sigma_{c}^{-1/2}P_{i}P_{i}^{\top}\Sigma_{c}^{-1/2}\mu$
$=\frac{1}{n_{c}}\mu^{\top}\Sigma_{c}^{-1/2}\left(\frac{1}{n_{c}}\sum_{\{i:A_{i}=0\}}P_{i}P_{i}^{\top}\right)\Sigma_{c}^{-1/2}\mu$
@@
where $d=\Sigma_{c}^{-1/2}\mu$ and $Z_{i}=P_{i}^{\top}d$ where $Z_{i}$’s are mean 0 and independent. Therefore, we have:
@@
$=\mathbb{E}\left[\mathbb{P}\left(\left|\sqrt{n}\sum_{\{i:A_{i}=0\}}\tau_{2,i}\rho_{i}\right|\geq t\mid\tilde{\mathcal{G}}_{n}\right)\mathds{1}_{\Omega_{n}}\right]$
@@
$=\frac{\sigma_{2}^{2}}{t^{2}}\mathbb{E}[\|\hat{\gamma}_{1}-\gamma_{1}\|_{2}^{2}]\mathbb{E}\left[\frac{1}{n_{c}}\sum_{\{i:A_{i}=0\}}Z_{i}^{2}\right]\hskip 36.135pt[\text{As $\hat{\gamma}_{1}$ is obtained from treatment observations}]$
$\leq\frac{\sigma_{2}^{2}}{t^{2}}\mathbb{E}[\|\hat{\gamma}_{1}-\gamma_{1}\|_{2}^{2}](\mu^{\top}\Sigma_{c}^{-1}\mu)$
[…]
@@ A.5 Calculation of Effects
Appendix B Sensitivity Analysis

### v3

@@ Appendix A Proofs
$T_{1}=\frac{1}{\sqrt{\sigma_{1}^{2}\|\tau_{1}\|_{2}^{2}+\sigma_{2}^{2}\|\tau_{2}\|_{2}^{2}\|\gamma_{1}\|_{2}^{2}}}\left(\sum_{\left\{i:A_{i}=1\right\}}\tau_{1,i}\epsilon_{i}+\sum_{\left\{i:A_{i}=0\right\}}\tau_{2,i}U_{i}^{\top}\gamma_{1}\right)\,$,
which follows by cancelling $\sqrt{n}$ from the definition of $\sigma_{n}$ in Equation (A.28). Since the noise is centered i.i.d. sub-Gaussian random variables (Assumption 5.6), we have conditional on $\mathcal{G}_{n}=\sigma(\mathcal{D}_{1},\mathbf{X},\mathbf{M}_{c},A_{1:n})$, $T_{1}$ has mean 0 and variance 1. Furthermore, it is easy to check that Lyapunov’s condition is also satisfied with $\delta=1$ as $\sqrt{n}\|\tau_{1}\|_{\infty}+\sqrt{n}\|\tau_{2}\|_{\infty}=o_{p}(1)$ and $\|\gamma_{1}\|_{2}$ is bounded. We define the individual terms as $T_{1,i}^{\prime}$ for $1\leq i\leq n_{t}+n_{c}$, where:
$T^{\prime}_{1,i}=\begin{cases}\tau_{1,i}\epsilon_{i},&\text{ if }A_{i}=1\\ \tau_{2,i}U_{i}^{\top}\gamma_{1},&\text{ if }A_{i}=0\,.\end{cases}$ (A.29)
Then, we have:
$\sum_{j=1}^{n}T^{\prime}_{1,j}=\sum_{\left\{i:A_{i}=1\right\}}\tau_{1,i}\epsilon_{i}+\sum_{\left\{i:A_{i}=0\right\}}\tau_{2,i}U_{i}^{\top}\gamma_{1}\,.$
As a consequence, we have:
$s_{n}^{2}=\operatorname{Var}\left(\sum_{j=1}^{n}T^{\prime}_{1,j}\mid\mathcal{G}_{n}\right)=\sigma_{1}^{2}\|\tau_{1}\|^{2}_{2}+\sigma^{2}_{2}\|\tau_{2}\|_{2}^{2}\|\gamma_{1}\|_{2}^{2}\,.$
We need to check that for $\delta=1$, the Lyapunov’s condition:
$\lim_{n\rightarrow\infty}\frac{1}{s_{n}^{2+\delta}}\sum_{j=1}^{n}\mathbb{E}[|T_{1,j}^{\prime}-0|^{2+\delta}\mid\mathcal{G}_{n}]=0\,$, (A.30)
is satisfied, then a sum of $\frac{T_{1,j}^{\prime}-0}{s_{n}}$ converges in distribution to $\mathcal{N}(0,1)$ as $n\rightarrow\infty$:
$\frac{1}{s_{n}}\sum_{i=1}^{n}(T_{1,j}^{\prime}-0)\mid\mathcal{G}_{n}\xrightarrow[]{\mathcal{D}}\mathcal{N}(0,1)\,$,
where $s_{n}=\sqrt{\sigma_{1}^{2}\|\tau_{1}\|_{2}^{2}+\sigma_{2}^{2}\|\tau_{2}\|_{2}^{2}\|\gamma_{1}\|_{2}^{2}}$. Therefore, we check that
$\sum_{j=1}^{n}\mathbb{E}[|T_{1,i}^{\prime}-0|^{3}\mid\mathcal{G}_{n}]$
$=\sum_{\{i:A_{i}=1\}}\mathbb{E}[|\tau_{1,i}\epsilon_{i}|^{3}\mid\mathcal{G}_{n}]+\sum_{\{i:A_{i}=0\}}\mathbb{E}[|\tau_{2,i}U_{i}^{\top}\gamma_{1}|^{3}\mid\mathcal{G}_{n}]$
$=\sum_{\{i:A_{i}=1\}}\tau_{1,i}^{3}\mathbb{E}[|\epsilon_{i}|^{3}\mid\mathcal{G}_{n}]+\sum_{\{i:A_{i}=0\}}\tau_{2,i}^{3}\mathbb{E}[|U_{i}^{\top}\gamma_{1}|^{3}\mid\mathcal{G}_{n}]$
$\leq C_{1}(v_{1}^{2}S_{1})^{3}n_{t}^{-2/3}\|\tau_{1}\|_{2}^{2}$
[…]
$=\mathcal{O}_{p}\left(n^{-2/3}(\|\tau_{1}\|^{2}_{2}+\|\tau_{2}\|^{2}_{2})\right)\hskip 36.135pt[n_{t}\sim n_{c}\sim n]$
for some universal constants $C_{1}$ and $C_{2}$. Next, we focus on the lower bound of $s_{n}^{3}$, we have:
$s_{n}^{3}$ $\geq\operatorname{max}\{(\sigma_{1}^{2}\|\tau_{1}\|^{2}_{2})^{3/2},(\sigma^{2}_{2}\|\tau_{2}\|_{2}^{2}\|\gamma_{1}\|_{2}^{2})^{3/2}\}\$
$\geq\frac{1}{2}(\sigma_{1}^{3}\|\tau_{1}\|^{3}_{2}+\sigma^{3}_{2}\|\tau_{2}\|_{2}^{3}\|\gamma_{1}\|_{2}^{3})$
[…]
$=\mathcal{O}_{p}\left(n^{-1/2}(\|\tau_{1}\|^{2}_{2}+\|\tau_{2}\|^{2}_{2})\right)\,$,
for some constant $C_{1},C_{2}>0$. Hence, we conclude that for $\delta=1$,
$\lim_{n\rightarrow\infty}\frac{1}{s_{n}^{2+\delta}}\sum_{i=1}^{n}\mathbb{E}[|T_{1,i}^{\prime}-0|^{2+\delta}\mid\mathcal{G}_{n}]=o_{p}(1)\,.$
This verifies the Lyapunov’s condition in Equation (A.30). Therefore, an application of the conditional central limit theorem (Bulinski, (2017)) yields:
@@
$\sum_{\{i:A_{i}=0\}}\tau_{2,i}^{*2}$ $=\sum_{\{i:A_{i}=0\}}\left(\frac{1}{n_{c}}\mu_{X}^{\top}\Sigma_{X,c}^{-1}(X_{c,i}-\mu_{X,c})\right)^{2}$
$=\frac{1}{n_{c}^{2}}\sum_{\{i:A_{i}=0\}}\left(\mu_{X}^{\top}\Sigma_{X,c}^{-1/2}P_{i}\right)^{2}$
$=\frac{1}{n_{c}^{2}}\sum_{\{i:A_{i}=0\}}\mu_{X}^{\top}\Sigma_{X,c}^{-1/2}P_{i}P_{i}^{\top}\Sigma_{X,c}^{-1/2}\mu_{X}$
$=\frac{1}{n_{c}}\mu_{X}^{\top}\Sigma_{X,c}^{-1/2}\left(\frac{1}{n_{c}}\sum_{\{i:A_{i}=0\}}P_{i}P_{i}^{\top}\right)\Sigma_{X,c}^{-1/2}\mu_{X}$
@@
where $d=\Sigma_{X,c}^{-1/2}\mu_{X}$ and $Z_{i}=P_{i}^{\top}d$ where $Z_{i}$’s are mean 0 and independent. Therefore, we have:
@@
$=\mathbb{E}\left[\left(\left|\sqrt{n}\sum_{\{i:A_{i}=0\}}\tau_{2,i}\rho_{i}\right|\geq t\mid\tilde{\mathcal{G}}_{n}\right)\mathds{1}_{\Omega_{n}}\right]$
@@
$=\frac{\sigma_{2}^{2}}{t^{2}}\mathbb{E}[\|\hat{\gamma}_{1}-\gamma_{1}\|_{2}^{2}]\mathbb{E}\left[\frac{1}{n_{c}}\sum_{\{i:A_{i}=0\}}Z_{i}^{2}\right]\hskip 7.22743pt[\text{As $\hat{\gamma}_{1}$ is obtained from treatment observations}]$
$\leq\frac{\sigma_{2}^{2}}{t^{2}}\mathbb{E}[\|\hat{\gamma}_{1}-\gamma_{1}\|_{2}^{2}](\mu_{X}^{\top}\Sigma_{X,c}^{-1}\mu_{X})$
$\leq V_{2}\frac{\sigma_{2}^{2}}{t^{2}}\mathbb{E}[\|\hat{\gamma}_{1}-\gamma_{1}\|_{2}^{2}]\,.$
@@ Appendix B Calculation of Effects
Appendix C Sensitivity Analysis
@@
Appendix D Phenotype Features
Feature Name Explanation
gender Patient’s gender (e.g., male or female).
age Age of the patient at diagnosis.
ABSOLUTE_Purity Tumor purity score estimated through molecular analysis.
number_pack_years_smoked Number of cigarette pack years smoked by the patient.
initial_weight Patient’s initial body weight at diagnosis.
radiation_therapy Indicates whether the patient underwent radiation therapy (Yes/No).
dlco_predictive_percent Predicted percentage of the diffusing capacity of the lungs for carbon monoxide.
intermediate_dimension Intermediate dimension of the tumor (e.g., size-related measurement).
longerest_dimension Longest dimension of the tumor.
karnofsky_performance_score Performance status score that quantifies a patient’s ability to perform daily activities.
anatomic_neoplasm_subdivision Anatomic location of the tumor in the lungs.
histological_type Histological classification of the tumor (e.g., adenocarcinoma, squamous cell carcinoma).
days_to_collection Number of days from diagnosis to the collection of samples.
new_tumor_event_ after_initial_treatment Indicates whether a new tumor event occurred post-treatment.
post_bronchodilator_ fev1_percent Predicted percentage of forced expiratory volume in one second post-bronchodilator use.
shortest_dimension Shortest dimension of the tumor.
pathologic_stage Clinical stage of the cancer (e.g., Stage I, Stage II, etc.).
sample_type Type of the sample (e.g., primary tumor, normal tissue).
targeted_molecular_therapy Indicates whether the patient received molecularly targeted therapy.
status Survival status of the patient (e.g., alive or deceased).